OpenStudy (anonymous):

Solve x2 + 4x – 12 = 0 by completing the square.

OpenStudy (anonymous):

x^2 + 4x +(4/2)^2 - (4/2)^2 – 12 = 0 x^2+4x+4-16=0 (x+2)^2=16 x+2=+-4 x=4-2=2 x=-4-2=-6

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw